OpenStudy (anonymous):

How many solutions does tan^2(x)-25=0 have on interval 0,2pi i can get to tan(x)=5 but what then.

jimthompson5910 (jim_thompson5910):

keep in mind that if x^2 = 25 then x = 5 or x = -5

jimthompson5910 (jim_thompson5910):

since (5)^2 = 25 and (-5)^2 = 25

jimthompson5910 (jim_thompson5910):

that means you should have these two equations tan(x) = 5 or tan(x) = -5

jimthompson5910 (jim_thompson5910):

you then apply the arctan function to both sides for each equation to solve for x

OpenStudy (anonymous):

ok i did, 1.373400766945015860861271926445 and -1.373400766945015860861271926445 However only one is in the interval, yet the other is within the range that i can add another pi to get 1.373400766945015860861271926445 and 4.5149934205348090993239153097245 and 1.7681918866447773776013714568345 and 4.909784540234570616064014840114 So there are 4 possible solutions, not two?

OpenStudy (anonymous):

Answer please?

jimthompson5910 (jim_thompson5910):

I count four solutions as well

jimthompson5910 (jim_thompson5910):

makes sense because each equation tan(x) = 5, tan(x) = -5 brings 2 solutions each